Name: Peter Zona

Content and Structure – The paper

1. Has a clearly stated topic or question that it will address – 10

2. Does address the topic or question – 5

3. Has a clearly articulated and logical structure:

a. First paragraph lays out the topic or question – 8 b. Later paragraphs have topic sentences and clearly identifiable main ideas – 10 c. Transitions between paragraphs are logical – 10

4. Accurately presents facts and viewpoints of sources – 10

5. Provides sufficient textual evidence to support the argument – 10

6. Is concise, clear, and not too repetitive – 10

7. Has a clearly identifiable conclusion – 10

8. Has a convincing argument that draws together the evidence presented – 10

Total: 93 (A-)

Style and English Usage – The paper

1. Is clear, direct, and appropriate in tone – 20

2. Is free of misspellings – 10

3. Is free of grammatical mistakes – 15

4. Avoids incomplete sentences – 10

5. Avoids run-on or overly complicated sentences – 10

6. Uses correct punctuation – 10

7. Is free of jargon, cliches, and misused words – 10

8. Cites quotations and/or references correctly – 10

Total: 100 (A)

Comments: Peter, This is a good improvement over the first draft. However, think that your conclusion about ethics (that an ethical sense is ingrained at birth) is questionable. It seems to me that each of us also forms our sense of ethics through the interactions we have with other people as we grow up. Our parents, teachers, priests, friends, and others all have a part in that. For that reason, I think that (as you said in your opening paragraph) there are times when we have to say that some things are wrong (even if they might be legal). I don’t know for sure, but I would be willing to bet that most people would not agree that doctors should be able to use their tissues without their consent in the way that Skloot is discussing on page 315. So whether or not that is legal, there is still the question whether it should be legal. I think that is the ultimate question Skloot is trying to pose here.